OpenStudy (anonymous):

Use an inequality to solve the problem. Be sure to show the inequality and all of your work. A flag maker is making a triangular flag, but he must keep the sign under 40 square feet. If the base of the sign is 10 feet, what possible values of the height of the triangular sign are allowed? please help me

OpenStudy (anonymous):

Do you know the formula for area of a triangle? Start with that

OpenStudy (anonymous):

No. I have no idea how to do this. I have been out of school since 1985 and never had to take algebra . now it's kicking my behind.

OpenStudy (anonymous):

Oh that's a long time. Well the formula is \[ A = \frac{b * h}{2} \] Where b is the lenght of the base of the triangle and h is the height of the triangle. So can you think what to do next now?

OpenStudy (anonymous):

PLUG IN THE NUMBERS?

OpenStudy (anonymous):

Yes and the equals sign must be replaced with > or <. And since A can be at most 40 what way do you think the > should face?

OpenStudy (anonymous):

I'M JUST AS DUMB AS A ONION. I HAVE NO IDEA HOW TO DO THIS STUFF. I APPRECIATE YOU HELPING ME! I HAVE NO BOOK AND IT'S HARD FOR ME TO JUST LOOK AT SOMETHING AND DON'T KNOW HOW TO DO IT

OpenStudy (anonymous):

Ok, so lets do it together. Lets plug in the numbers, that is the ones that we have. So area must be 40 and base must be 10. \[ 40 = \frac{10*b}{2} \] All ok here?

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

i made a small mistake. the formula is now \[ 40 feet^2 = \frac{10 feet * h}{2} \] (we pluged in A and b not A and h.. nevermind that) So now the ">" thingy. On the right side of the equals sign we have area that we get if we plug in some value for h and on the left is how big the flag CAN be. The right way to put it is \[ 40 feet^2 > \frac{10 feet * h}{2} \] So that 40 square feet is the limiting value of the area. The flag can only be as big as that.

OpenStudy (anonymous):

Thank you for being patient with and helping me. I really appreciate it. I will have to find a tutor. being 45 and trying to go back to school is driving me insane.

OpenStudy (anonymous):

No problem, that's what this place is for. Do you know the answer now?

OpenStudy (anonymous):

isnt that the answer?

OpenStudy (anonymous):

see i told you i cant do this. lol

OpenStudy (anonymous):

Sry. No not the solution yet. I thin you must show what h can be most. You must multiplay both sides by 2 and devied both by 10 feet to get \[ 8 feet > h \] now you know that h can be at most 8 feet and no more in order not to exceed the 40 square feet condition.

OpenStudy (anonymous):

THANKS